# Toward Robust, Reproducible, and Widely Accessible Intracranial Language Brain-Computer Interfaces: A Comprehensive Review of Neural Mechanisms, Hardware, Algorithms, Evaluation, Clinical Pathways and Future Directions

### 03 Mar 2026 v1

@@ 3.2 Quantitative Trade-offs in Resolution, Signal Quality, and Clinical Feasibility
In 2023, Willett et al. [128] conducted a comprehensive investigation on intracranial MEAs, providing key insights into their performance metrics and comparisons with other intracranial recording modalities. In terms of spatial sampling density, the 3.2 x 3.2 mm² MEAs used in their study contained 128 channels each, exhibiting rich and spatially intermixed tuning to speech articulators, which allows for precise decoding from a small cortical area [128]. For signal quality, threshold crossing rates with root-mean-square (RMS) thresholds ranging from -4.5 to -3.5 effectively detected strong spiking activity well above the noise floor; notably, combining spike band power and threshold crossings was found to improve RNN decoding performance. Regarding coverage depth, the arrays were implanted intracortically, targeting superficial cortical layers in area 6v, and while their cortical coverage was narrow, this was offset by the high neural information content per channel. In terms of decoding performance, intracortical MEAs paired with RNN decoders achieved a high decoding speed of 62 words per minute, which is 3.4 times faster than previous BCI speech decoding approaches [128]. WERs were reported as 9.1% on a 50-word vocabulary and 23.8% on a 125,000-word vocabulary, with error rates decreasing log-linearly as the electrode count increased—specifically, doubling the number of electrodes nearly halved the error rate [128].
When considering task-modality matching, high-resolution MEAs are recommended for precise speech reconstruction at the phoneme and word levels, particularly when the goal is rapid and unconstrained communication with large vocabularies. In contrast, the study predicts that SEEG or macro-ECoG may be more suitable for clinical communication interfaces with simpler vocabularies, while MEAs are best suited for supporting high-fidelity speech synthesis or restoration. Comparisons to other modalities highlighted that ECoG grids reportedly show broader somatotopy but lower spatial resolution than MEAs. However, MEAs have notable constraints: they require surgical implantation, have limited spatial coverage, and their long-term stability and clinical viability need further validation. Additionally, training data requirements are high, with 260–440 sentences needed per day, though decoding still performs reasonably with reduced retraining [128]. The study also noted that Brodmann area 44, when sampled with intracortical electrodes, contained little to no useful speech-related information, emphasizing the importance of precise targeting for optimal performance.
@@ 4 Algorithmic Advances in Generalization and Interpretability
In the hardware–software co-design view of Fig. 2, algorithmic choices determine how effectively recorded neural information is converted into robust linguistic outputs under fixed recording constraints. Beyond subject-specific laboratory demonstrations, clinically viable language BCIs require decoders that generalize across participants, heterogeneous electrode layouts, sessions, and speech modes, while remaining interpretable enough to support mechanism validation and user-centered safeguards. This section therefore reviews recent advances spanning topology-agnostic sequence/transformer architectures, biologically grounded articulatory intermediates, language-prior assistance, dual-path/multi-modal decoding frameworks, and interpretable models for linking performance gains back to neural mechanisms.
@@ 4.1 Cross-Subject Generalization via Sequence and Transformer Architectures
Cross-subject generalization in language BCIs has been significantly advanced through the application of sequence and transformer architectures, which address the challenges of variable electrode placements and sparse cortical coverage. Sequence-to-sequence models have demonstrated robust performance in decoding speech-related neural signals across participants. In 2025, Singh et al. [111] employed a sequence-to-sequence model with LSTM layers to decode variable-length phonemic sequences, developing a cross-subject transfer learning framework that isolates shared latent manifolds. Their group-derived decoder outperformed models trained on individual data alone, with transfer learning involving pre-training a core RNN encoder and affine layers on a single subject, then freezing these components while fine-tuning a subject-specific 1D convolutional layer to accommodate variable electrode configurations; this approach resulted in no significant difference in PER when transferred to unseen participants. Similarly, in 2020, Makin et al. [80] implemented a sequence-to-sequence encoder-decoder architecture with RNN-LSTM cells to decode ECoG signals into written sentences, achieving mean WERs as low as 3% for individual participants. Transfer learning further improved cross-subject generalization, with networks pretrained on one participant and fine-tuned on limited data from another reducing WER by up to 36%, even with ECoG arrays in different hemispheres. For SEEG data, in 2024, Wu et al. [132] showed that both RNN-based and transformer-based sequence-to-sequence models outperformed linear regression baselines in mel-spectrogram reconstruction accuracy, with comparable results achievable using only a few electrodes, highlighting the robustness of these architectures to electrode sparsity.
@@
Transformer-based decoders have further enhanced cross-subject generalization by leveraging anatomical information and shared neural representations. In 2023, Défossez et al. [30] introduced a shared deep convolutional architecture with a participant-specific layer, trained on non-invasive MEG/EEG data using contrastive learning aligned to pretrained speech representations (wav2vec 2.0). This group-derived model outperformed individual-specific decoders, achieving up to 70.7% top-10 segment-level accuracy in MEG datasets and generalizing to unseen participants and vocabulary. A notable advancement is the SwinTW transformer, proposed by 2024, Chen et al. [18], which uniquely handles arbitrarily positioned electrodes by leveraging their 3D anatomical locations (MNI coordinates and brain region embeddings) instead of 2D grid indices. SwinTW generates tokens from each electrode individually and applies temporal window attention, enabling multi-subject training without subject-specific layers. Multi-subject SwinTW models achieved decoding performance comparable to subject-specific models on training participants (e.g., PCC of 0.837 vs. 0.831) and generalized to unseen participants with an average PCC of 0.765 in leave-one-out cross-validation, outperforming grid-based models like ResNet and 3D Swin Transformer. This is why topology-agnostic transformer families are placed near the high-generalization region in Fig. 5: their architectural inductive bias is designed around anatomical variability rather than fixed electrode grids, which directly targets the central barrier to cross-subject transfer.
@@ 4.5 Interpretable Decoding Models for Neural Mechanism Validation
Interpretable decoding models are critical for validating neural mechanisms underlying speech by establishing links between neural activity and articulatory or linguistic components, thereby advancing both scientific understanding and clinical BCI optimization. In 2025, Singh et al. [111] developed a sequence-to-sequence (Seq2Seq) model that integrates temporal convolutional networks, recurrent neural networks, and a linear readout layer to isolate phoneme identity probabilities, enabling interpretable mapping from neural data to phonemic sequences. This model outperformed linear baselines in phoneme prediction across patients and, through REO analysis, identified critical roles of ventral sensorimotor cortex, posterior superior temporal gyrus, and superior temporal sulcus in phonological processing—establishing region-to-phoneme mappings aligned with known neuroanatomical contributions to speech. Transfer learning with a shared recurrent layer further encoded subject-invariant articulatory information, producing latent feature embeddings that facilitated decoding in patients with variable electrode coverage and validated distributed neural mechanisms underlying speech. The use of PER provided an interpretable metric to quantify decoding precision at the phonemic level, bridging neural activity to articulatory and linguistic units.
Linear models have served as interpretable baselines for validating neural patterns, complementing more complex architectures. In 2022, Luo et al. [79] noted that phoneme decoding studies using linear classifiers achieved accuracies ranging from  20% to over 70% with intracranial recordings, directly mapping neural features to articulatory or linguistic components. Beyond linear approaches, deep learning models such as convolutional neural networks (CNNs) mapped high gamma activity to speech features like mel-spectrograms, while recurrent neural networks and encoder-decoder frameworks improved word and sentence decoding accuracy. Speech synthesis approaches, which first map neural activity to intermediary representations (e.g., spectrotemporal features or vocal tract trajectories) before acoustic reconstruction via vocoders like WaveNet, enhance interpretability by elucidating neural contributions to speech production mechanisms and supporting clinical BCI optimization.
@@ 5.2 Limitations of Conventional Metrics and Role of Perceptual Evaluation
Conventional objective metrics in speech neuroprosthesis evaluation, such as correlation coefficient (CC) and mel cepstral distortion (MCD), suffer from significant limitations and inconsistent usage, hindering meaningful cross-task comparisons. Correlation coefficient, widely relied upon in many studies, is sensitive to analysis parameters like window size and focuses on co-variation rather than absolute signal differences, failing to capture perceptual intelligibility or naturalness. Similarly, MCD, which operates on mel-frequency cepstral coefficients (MFCCs) and is preferred over correlation-based metrics for weighting perceptually relevant features, ignores temporal alignment and can unfairly penalize perceptually good reconstructions with constant spectral offsets, thus not fully reflecting speech intelligibility. The inconsistent application of these metrics, coupled with variations in parameters and experimental conditions, leads to non-comparable evaluation results across different studies and datasets.
@@ 6.5 Multi-Modal BCI Integration for Comprehensive Communication
Combining speech decoding with cursor control and other BCI modalities addresses critical limitations of single-modal systems, significantly enhancing real-world utility by leveraging the complementary strengths of each modality. In Fig. 7, this corresponds to the multimodal utility gate required for broader clinical communication tasks. Traditional non-language BCIs, such as those using P300, motor imagery, or steady-state visually evoked potential, cannot match the speed and flexibility of spoken communication. In contrast, language BCIs offer a more efficient and expressive channel by enabling real-time synthesis of speech, while integrating cursor control provides additional means for precise computer interaction or device manipulation. This multi-modal approach not only expands the range of communication and control capabilities but also accommodates varied user needs in dynamic environments.
@@
Evidence for shared cortical substrates supporting multi-modal BCI functions is increasingly robust. Studies have demonstrated that overlapping neural regions can support both speech decoding and motor control. For instance, an individual with tetraplegia successfully used ECoG signals recorded from the left sensorimotor cortex to operate 3D cursor movement by associating attempted movements (e.g., thumb, elbow, wrist flexion/extension) with cursor directions, while the same neural signals also supported speech-related attempted movements [126]. Similarly, recent research has shown that speech can be decoded from microelectrode Utah arrays implanted in dorsal motor areas, indicating that these regions might subserve multiple BCI modalities [79]. High-density recordings from peri-Sylvian cortices, including vSMC and STG, have further highlighted that these areas—associated with speech production and auditory feedback—can support accurate decoding of complex speech sequences, suggesting their potential to serve as shared substrates for speech and other control functions like cursor movement [80].
@@
The modular nature of BCI systems facilitates multi-modal integration. A typical synthesis-based language BCI comprises stages of neural signal recording, feature extraction, speech decoding, and audio synthesis, a framework that can be extended to incorporate additional modalities such as cursor control by integrating parallel decoding pathways. Advanced computational approaches, such as transfer learning across participants and sentence sets, enhance adaptability; networks pretrained on one task (e.g., speech decoding) can improve performance on another (e.g., cursor control), reflecting multi-task learning capacities. Furthermore, decoding lexical semantic information—such as the meaning of words—from neural activity, as demonstrated using SEEG recordings with 21% accuracy (vs. 10% chance) during spontaneous conversation, offers opportunities to combine semantic decoding with phoneme-focused speech models, potentially producing smoother and more accurate communication [99].
@@
The translational benefits of multi-modal functionality are substantial. First, shared cortical recording sites reduce the need for extensive implantations, improving safety and feasibility for clinical translation. The safety and stability of ECoG implants in individuals with late-stage ALS, which maintain high performance over long periods, further support the clinical viability of multi-modal systems. Second, multi-modal BCIs address real-world usability by providing redundant or complementary communication channels; for example, speech synthesis enables rapid expressive communication, while cursor control offers precise manipulation for tasks requiring fine motor control. Finally, integrating diverse modalities—such as speech, cursor movement, and semantic decoding—enhances overall system robustness and user satisfaction, critical factors for successful clinical adoption. Collectively, these advances position multi-modal BCI integration as a key pathway toward comprehensive communication restoration for individuals with speech impairments.
@@ 6.6 Clinical Trial Design for Speech Neuroprosthesis Validation
Participant selection criteria for speech neuroprosthesis clinical trials must balance clinical relevance and feasibility, focusing on individuals with speech impairment while ensuring safety and meaningful outcomes. This subsection addresses the regulatory validation gate in Fig. 7, where study design determines whether promising prototypes can generate approval-relevant evidence. For example, a study included a person with bulbar-onset ALS who could no longer speak intelligibly but retained limited orofacial movement and vocalization ability, illustrating the targeting of participants with residual motor function relevant to speech [128]. Similarly, the BrainGate2 pilot clinical trial enrolled participants with tetraplegia and residual speech ability to safely explore neural decoding performance [129], while another BrainGate2 participant with severe motor impairment (tetraplegic with dysarthria) demonstrated feasibility in a highly impaired population [110]. Surveys of potential users, including those with ALS and spinal cord injury, highlight that willingness to undergo implantation depends on the system meeting priorities like restoring communication with sufficient accuracy (>90%) [114].

### 16 Mar 2026 v2

@@ 3.2 Quantitative Trade-offs in Resolution, Signal Quality, and Clinical Feasibility
Willett et al. (2023) [130] conducted a comprehensive investigation on intracranial MEAs, providing key insights into their performance metrics and comparisons with other intracranial recording modalities. In terms of spatial sampling density, the 3.2 x 3.2 mm² MEAs used in their study contained 128 channels each, exhibiting rich and spatially intermixed tuning to speech articulators, which allows for precise decoding from a small cortical area [130]. For signal quality, threshold crossing rates with root-mean-square (RMS) thresholds ranging from -4.5 to -3.5 effectively detected strong spiking activity well above the noise floor. Notably, combining spike band power with threshold crossings was found to improve RNN decoding performance. Regarding coverage depth, the arrays were implanted intracortically, targeting superficial cortical layers in area 6v. Although their cortical coverage was narrow, this was offset by the high neural information content per channel. In terms of decoding performance, intracortical MEAs paired with RNN decoders achieved a high decoding speed of 62 words per minute, which is 3.4 times faster than previous BCI speech decoding approaches [130]. WERs were reported as 9.1% on a 50-word vocabulary and 23.8% on a 125,000-word vocabulary, with error rates decreasing log-linearly as the electrode count increased. Specifically, doubling the number of electrodes nearly halved the error rate [130].
In considering the matching of task to recording modality, high-resolution MEAs are recommended for precise speech reconstruction at the phoneme and word levels, particularly when the goal is rapid and unconstrained communication with large vocabularies. In contrast, the study predicts that SEEG or macro-ECoG may be more suitable for clinical communication interfaces with simpler vocabularies, while MEAs are best suited for supporting high-fidelity speech synthesis or restoration. Comparisons to other modalities highlighted that ECoG grids reportedly show broader somatotopy but lower spatial resolution than MEAs. However, MEAs have notable constraints: they require surgical implantation, have limited spatial coverage, and their long-term stability and clinical viability need further validation. Additionally, training data requirements are high, with 260–440 sentences needed per day, though decoding still performs reasonably with reduced retraining [130]. The study also noted that Brodmann area 44, when sampled with intracortical electrodes, contained little to no useful speech-related information, emphasizing the importance of precise targeting for optimal performance.
@@ 4 Algorithmic Advances in Generalization and Interpretability
In the hardware–software co-design view of Fig. 2, algorithmic choices determine how effectively recorded neural information is converted into robust linguistic outputs under fixed recording constraints. Beyond subject-specific laboratory demonstrations, clinically viable language BCIs require decoders that generalize across participants, heterogeneous electrode layouts, sessions, and speech modes, while remaining interpretable enough to support mechanism validation and user-centered safeguards. This section reviews recent advances spanning topology-agnostic sequence and transformer architectures, biologically grounded articulatory intermediates, language-prior assistance, dual-path and multi-modal decoding frameworks, and interpretable models for linking performance gains back to neural mechanisms.
@@ 4.1 Cross-Subject Generalization via Sequence and Transformer Architectures
Cross-subject generalization in language BCIs has been significantly advanced through the application of sequence and transformer architectures, which address the challenges of variable electrode placements and sparse cortical coverage. Sequence-to-sequence models have demonstrated robust performance in decoding speech-related neural signals across participants. Singh et al. (2025) [113] employed a sequence-to-sequence model with LSTM layers to decode variable-length phonemic sequences, developing a cross-subject transfer learning framework that isolates shared latent manifolds. Their group-derived decoder outperformed models trained on individual data alone, with transfer learning involving pre-training a core RNN encoder and affine layers on a single subject, then freezing these components while fine-tuning a subject-specific 1D convolutional layer to accommodate variable electrode configurations; this approach resulted in no significant difference in PER when transferred to unseen participants. Similarly, Makin et al. (2020) [82] implemented a sequence-to-sequence encoder-decoder architecture with RNN-LSTM cells to decode ECoG signals into written sentences, achieving mean WERs as low as 3% for individual participants. Transfer learning further improved cross-subject generalization, with networks pretrained on one participant and fine-tuned on limited data from another reducing WER by up to 36%, even with ECoG arrays in different hemispheres. For SEEG data, Wu et al. (2024) [134] showed that both RNN-based and transformer-based sequence-to-sequence models outperformed linear regression baselines in mel-spectrogram reconstruction accuracy, with comparable results achievable using only a few electrodes, highlighting the robustness of these architectures to electrode sparsity.
@@
Transformer-based decoders have further enhanced cross-subject generalization by leveraging anatomical information and shared neural representations. Défossez et al. (2023) [31] introduced a shared deep convolutional architecture with a participant-specific layer, trained on non-invasive MEG/EEG data using contrastive learning aligned to pretrained speech representations (wav2vec 2.0). This group-derived model outperformed individual-specific decoders, achieving up to 70.7% top-10 segment-level accuracy in MEG datasets and generalizing to unseen participants and vocabulary. A notable advancement is the SwinTW transformer, proposed by Chen et al. (2024) [19], which uniquely handles arbitrarily positioned electrodes by leveraging their 3D anatomical locations (MNI coordinates and brain region embeddings) instead of 2D grid indices. SwinTW generates tokens from each electrode individually and applies temporal window attention, enabling multi-subject training without subject-specific layers. Multi-subject SwinTW models achieved decoding performance comparable to subject-specific models on training participants (e.g., PCC of 0.837 vs. 0.831) and generalized to unseen participants with an average PCC of 0.765 in leave-one-out cross-validation, outperforming grid-based models like ResNet and 3D Swin Transformer. This is why topology-agnostic transformer families are placed near the high-generalization region in Fig. 5: their architectural inductive bias is designed around anatomical variability rather than fixed electrode grids, which directly targets the central barrier to cross-subject transfer.
@@ 4.5 Interpretable Decoding Models for Neural Mechanism Validation
Interpretable decoding models are critical for validating neural mechanisms underlying speech by establishing links between neural activity and articulatory or linguistic components, thereby advancing both scientific understanding and clinical BCI optimization. Singh et al. (2025) [113] developed a sequence-to-sequence (Seq2Seq) model that integrates temporal convolutional networks, recurrent neural networks, and a linear readout layer to isolate phoneme identity probabilities, enabling interpretable mapping from neural data to phonemic sequences. This model outperformed linear baselines in phoneme prediction across patients and, through REO analysis, identified critical roles of ventral sensorimotor cortex, posterior superior temporal gyrus, and superior temporal sulcus in phonological processing—establishing region-to-phoneme mappings aligned with known neuroanatomical contributions to speech. Transfer learning with a shared recurrent layer further encoded subject-invariant articulatory information, producing latent feature embeddings that facilitated decoding in patients with variable electrode coverage and validated distributed neural mechanisms underlying speech. The use of PER provided an interpretable metric to quantify decoding precision at the phonemic level, bridging neural activity to articulatory and linguistic units.
Linear models have served as interpretable baselines for validating neural patterns, complementing more complex architectures. Luo et al. (2022) [81] noted that phoneme decoding studies using linear classifiers achieved accuracies ranging from  20% to over 70% with intracranial recordings, directly mapping neural features to articulatory or linguistic components. Beyond linear approaches, deep learning models such as convolutional neural networks (CNNs) mapped high gamma activity to speech features like mel-spectrograms, while recurrent neural networks and encoder-decoder frameworks improved word and sentence decoding accuracy. Speech synthesis approaches, which first map neural activity to intermediary representations (e.g., spectrotemporal features or vocal tract trajectories) before acoustic reconstruction via vocoders like WaveNet, enhance interpretability by elucidating neural contributions to speech production mechanisms and supporting clinical BCI optimization.
@@ 5.2 Limitations of Conventional Metrics and Role of Perceptual Evaluation
Conventional objective metrics in speech neuroprosthesis evaluation, such as correlation coefficient (CC) and mel cepstral distortion (MCD), suffer from significant limitations and inconsistent usage, hindering meaningful cross-task comparisons [133]. Correlation coefficient, widely relied upon in many studies, is sensitive to analysis parameters like window size and focuses on co-variation rather than absolute signal differences, failing to capture perceptual intelligibility or naturalness. Similarly, MCD, which operates on mel-frequency cepstral coefficients (MFCCs) and is preferred over correlation-based metrics for weighting perceptually relevant features, ignores temporal alignment and can unfairly penalize perceptually good reconstructions with constant spectral offsets, thus not fully reflecting speech intelligibility. The inconsistent application of these metrics, coupled with variations in parameters and experimental conditions, leads to non-comparable evaluation results across different studies and datasets.
@@ 6.5 Multi-Modal BCI Integration for Comprehensive Communication
Combining speech decoding with cursor control and other BCI modalities addresses critical limitations of single-modal systems by leveraging the complementary strengths of each approach, thereby enhancing real-world utility. In Fig. 7, this corresponds to the multimodal utility gate required for broader clinical communication tasks. Traditional non-language BCIs, such as those using P300, motor imagery, or steady-state visually evoked potential, cannot match the speed and flexibility of spoken communication. Language BCIs, by contrast, offer a more efficient and expressive channel through real-time synthesis, while the integration of cursor control provides an additional means of precise computer interaction or device manipulation. This multi-modal approach not only expands the range of communication and control capabilities but also accommodates varied user needs in dynamic environments.
@@
Evidence for shared cortical substrates supporting multi-modal BCI functions is increasingly robust, with studies demonstrating that overlapping neural regions can support both speech decoding and motor control. For instance, an individual with tetraplegia successfully used ECoG signals recorded from the left sensorimotor cortex to operate 3D cursor movement by associating attempted movements (e.g., thumb, elbow, wrist flexion/extension) with cursor directions, while the same neural signals simultaneously supported speech movements [128]. Similarly, recent research has shown that speech can be decoded from microelectrode Utah arrays implanted in dorsal motor areas, indicating that these regions may subserve multiple BCI modalities [81]. High-density recordings from peri-Sylvian cortices, including vSMC and STG, have further revealed that these areas, which are associated with speech production and auditory feedback, can support accurate decoding of complex speech sequences. This finding points to their potential as shared substrates for speech and other control functions, such as cursor movement [82].
@@
The modular nature of BCI systems facilitates multi-modal integration. A typical synthesis-based language BCI comprises stages of neural signal recording, feature extraction, speech decoding, and audio synthesis, a framework that can be extended to incorporate additional modalities such as cursor control by integrating parallel decoding pathways. Advanced computational approaches, such as transfer learning across participants and sentence sets, enhance adaptability by enabling networks pretrained on one task (e.g., speech decoding) to improve performance on another (e.g., cursor control), reflecting multi-task learning capacities. Furthermore, decoding lexical semantic information—such as word meanings—from neural activity, as demonstrated using SEEG recordings with 21% accuracy (vs. 10% chance) during spontaneous conversation, offers opportunities to combine semantic decoding with phoneme-focused speech models, potentially producing smoother and more accurate communication [101].
@@
The translational benefits of multi-modal functionality are substantial. First, shared cortical recording sites reduce the need for extensive implantations, enhancing safety and feasibility for clinical translation. This is supported by evidence that ECoG implants in individuals with late-stage ALS remain stable and maintain high performance over long periods, further reinforcing the clinical viability of multi-modal systems. Second, multi-modal BCIs address real-world usability by providing redundant or complementary communication channels: speech synthesis enables rapid expressive communication, while cursor control offers precise manipulation for tasks requiring fine motor coordination. Finally, integrating diverse modalities—such as speech, cursor movement, and semantic decoding—enhances overall system robustness and user satisfaction, both critical for successful clinical adoption. Collectively, these advances position multi-modal BCI integration as a key pathway toward comprehensive communication restoration for individuals with speech impairments.
@@ 6.6 Clinical Trial Design for Speech Neuroprosthesis Validation
Participant selection criteria for speech neuroprosthesis clinical trials must balance clinical relevance and feasibility, focusing on individuals with speech impairment while ensuring safety and meaningful outcomes. This subsection addresses the regulatory validation gate in Fig. 7, where study design determines whether promising prototypes can generate approval-relevant evidence. For example, a study included a person with bulbar-onset ALS who retained limited orofacial movement and vocalization ability despite being unable to speak intelligibly, highlighting the importance of recruiting participants with residual speech-related motor function [130]. Similarly, the BrainGate2 pilot clinical trial enrolled participants with tetraplegia and residual speech ability to safely explore neural decoding performance [131], while another BrainGate2 participant with severe motor impairment (tetraplegic with dysarthria) demonstrated feasibility in a highly impaired population [112]. Surveys of potential users, including those with ALS and spinal cord injury, highlight that willingness to undergo implantation depends on the system meeting priorities like restoring communication with sufficient accuracy (>90%) [116].